# Chinese Academy of Sciences, Beijing 100190, P. R. China

### v1

@@ 4 Critical collapse
Let us assume that the perturbed region is spherically symmetric and has an initial radius $R(t)$ larger than the Hubble scale. The locally overdense region can then constitute a spatially closed universe with a metric given by Carr:1975qj ; Yokoyama:1995ex
$ds^{2}=-dt^{\prime 2}+R^{2}\left(t^{\prime}\right)\left[\frac{dr^{2}}{1-\kappa r^{2}}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right)\right],$ (18)
@@
$\rho=\frac{3}{8\pi G}\left(H^{2}+\frac{\kappa}{R^{2}}\right).$ (19)
The initial expansion rates of the background and the perturbed region can be set to be the same by choosing the coordinates Harrison:1969fb ; Yokoyama:1995ex so that the initial density contrast is given by
$\delta_{i}=\frac{\rho_{i}-\bar{\rho}_{i}}{\bar{\rho}_{i}}=\frac{3\kappa}{8\pi G\bar{\rho}_{i}R_{i}^{2}}=\frac{\kappa}{H_{i}^{2}R_{i}^{2}},$ (20)
@@
$\rho=\rho_{i}\left(\frac{R}{R_{i}}\right)^{-3(1+\omega)},$ (22)
@@
$\frac{\rho_{c}}{\rho_{i}}=\frac{\kappa}{R_{i}^{2}H_{i}^{2}+\kappa}\frac{R_{i}^{2}}{R_{c}^{2}}=\frac{\delta_{i}}{1+\delta_{i}}\frac{R_{i}^{2}}{R_{c}^{2}}.$ (23)
Together with Eqs. (22) and (23), we obtain the critical scale factor that the overdense region stops expanding
$\frac{R_{c}}{R_{i}}=\left(\frac{1+\delta_{i}}{\delta_{i}}\right)^{\frac{1}{1+3\omega}}.$ (24)
@@
$\frac{t_{c}}{t_{i}}=\left(\frac{1+\delta_{i}}{\delta_{i}}\right)^{\frac{3(1+\omega)}{2(1+3\omega)}}.$ (25)
The overdense region could contrast further against the pressure gradient if the self-gravitation $\Phi$ of the mass within the critical radius $R_{c}$ is larger than the internal energy $U$ of the overdense region, which gives Carr:1974nx
$\Phi+U\sim-G\rho^{2}R_{c}^{5}+\omega\rho R_{c}^{3}\gtrsim 0.$ (26)
@@
$R_{c}\gtrsim R_{J}\simeq c_{s}t_{c},$ (27)
where $R_{J}$ is the Jeans scale and $c_{s}\simeq\sqrt{\omega}$. With Eqs. (24), (25), and (27), one finds the critical density criterion for PBH formation in the radiation-dominated universe:
$\delta_{i}\gtrsim\delta_{c}\simeq c_{s}^{2}\simeq\frac{1}{3}.$ (28)
@@
$\frac{1}{3}\lesssim\delta_{i}\lesssim 1.$ (29)
Since the density perturbations are induced by the number density fluctuations of the domain walls, the scale of the density perturbation should have a characteristic size as that of the walls and the evolution of the perturbation should also follow the evolution of the walls. As shown in section 2.2, the domain wall in the scaling regime with density $\rho_{w}\simeq\sigma_{w}/t$ has a horizon-size characteristic length $\overline{L}\simeq t$. The number density fluctuations of the wall during radiation domination would generate horizon-size density fluctuations, whose average amplitude scales with time as
[…]
A rough estimate using $\delta\sim\rho_{w}/\rho_{\rm rad}\sim G\sigma_{w}t$ would lead to the same conclusion. Domain walls are continuous sources for inducing density perturbations. At a given time, the density fluctuations are randomly generated by the number density fluctuation. Equation (30) indicates that the probability for the production of a large-density contrast grows with the increase of cosmic time. Those perturbations with amplitudes that exceed the critical value $\delta_{c}$ might undergo critical collapse and lead to PBH formation.
We note in passing that a recent study Flores:2024lng argued that the critical density criterion could not be applied to horizon-size density perturbations in the inflationary scenarios since the over-density would be generated by the curvature perturbation, without which the horizon-size patch had only expansionary solutions. We note that the condition of having a superhorizon-size perturbation is the result of inflation, rather than the requirement for the critical collapse. Obviously, the domain wall scenario we consider in this work is not inflation-based and does not suffer from such a problem. In fact, the initial estimate by Hawking and Carr Carr:1974nx required merely a horizon-size perturbation for the critical collapse.
@@ 5.1 PBH relic abundance
$\beta(t_{f})=\frac{\rho_{\mathrm{PBH}}(t_{f})}{\rho_{\rm tot}(t_{f})}=\Omega_{\mathrm{PBH}}\left(\frac{H_{0}}{H_{f}}\right)^{2}\left(\frac{a_{0}}{a_{f}}\right)^{3},$ (31)
where $\Omega_{\mathrm{PBH}}\equiv\rho_{\mathrm{PBH}}\left(t_{0}\right)/\rho_{\text{crit}}$ (with $\rho_{\rm crit}=3H_{0}^{2}/(8\pi G)$) is the density of the PBH relic abundance at present and 00 denotes the present time. In the second equality of Eq. (31) we have used $\rho_{\rm PBH}(t_{f})=\rho_{\rm PBH}(t_{0})(a_{0}/a_{f})^{3}$. It is commonly related to the dark matter relic abundance by $\Omega_{\mathrm{PBH}}=f_{\mathrm{PBH}}\Omega_{\mathrm{CDM}}$, with the fraction of dark matter $f_{\rm PBH}$ in the form of PBHs.
It is expected that the PBHs formed from the Hubble horizon (re-)entry of the large density perturbations could have masses at the same order as the associated horizon mass $M_{H}\simeq 2\times 10^{15}(t/1s)M_{\odot}$. To further estimate the PBH production, let’s take the naive relation between the PBH mass $M$ and the horizon mass from critical phenomena Sasaki:2018dmp ; Carr:2020gox
$M=\gamma_{p}M_{H},$ (32)
@@
$M_{H}=\frac{4\pi}{3}\rho_{\rm rad}r_{H}^{3}=\frac{1}{2GH}=\frac{3\sqrt{5}M_{\rm pl}^{3}}{4\pi^{3/2}g_{*}^{1/2}(T)T^{2}}.$ (34)
Taking advantage of the Press-Schechter formalism, the PBH energy fraction can be expressed as Niemeyer:1997mt
[…]
where we consider a Gaussian probability distribution (16) for the density contrast. Note that we use an upper limit of infinity for the integration and confirm that it is not different from the integration result with an upper limit of $\delta=1$ from the requirement of avoiding separate universes Carr:1974nx ; Yokoyama:1995ex . A recent work Kopp:2010sh argued that this upper limit was a result of the geometry of the selected slicing rather than as a consequence of avoiding separate universes, which instead led to a divergence of density fluctuations on the superhorizon scale. With the PBH mass given by Eq. (32), the fraction of PBHs at the formation time is
$\beta(M)=\gamma_{p}{\rm erfc}\left(\frac{\delta_{c}}{\sqrt{2}\sigma}\right),$ (36)
where ${\rm erfc}(x)$ is the complementary error function (see appendix D for more details). The critical density contrast $\delta_{c}$ varies with the equation of state $\omega$. For radiation phase $\delta_{c}=0.453$. We determine $\delta_{c}$ with the $\delta_{c}-\omega$ relation revealed by the simulations in Ref. Musco:2012au (see section 6).
A more precise PBH mass formation was obtained from the numerical computations on the collapse of a relativistic fluid, which indicated a simple scaling law for the PBH mass distribution function near the critical threshold Choptuik:1992jv ; Evans:1994pj ; Koike:1995jm
$M=CM_{H}(\delta-\delta_{c})^{b},$ (37)
where the constant parameter $C=3.3$ and $b=0.36$ for the formation of PBH during radiation domination Byrnes:2018clq . Note that these parameters may be affected by the phase of cosmic matter. Then the PBH energy fraction is given by
[…]
In this work, we use the PBH mass (32) by default, unless otherwise stated.
@@
$s_{f}a_{f}^{3}=s_{0}a_{0}^{3},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm with}\leavevmode\nobreak\ \leavevmode\nobreak\ s(T)=\frac{2\pi^{2}}{45}g_{*,s}T^{3},$ (39)
where the current universe entropy density $s_{0}=2891.2\leavevmode\nobreak\ \rm cm^{-3}$. Together with Eqs. (31)-(33) we obtain the the current PBH relic density abundance
$\Omega_{\mathrm{PBH}}$ $=\frac{2\pi G\beta(M)s_{0}g_{*}(T)T}{H_{0}^{2}g_{*,s}(T)}=(3\sqrt{5\pi})^{1/2}\frac{\beta(M)s_{0}}{H_{0}^{2}(M_{\rm pl}M_{H})^{1/2}}\frac{g_{*}^{3/4}(T)}{g_{*,s}(T)}$ (40)
@@
where the Hubble expansion rate is $H_{0}=100h\leavevmode\nobreak\ \mathrm{km\leavevmode\nobreak\ s}^{-1}\leavevmode\nobreak\ \mathrm{Mpc}^{-1}$ with $h=0.68$ from the Planck observation Planck:2018vyg . With the relation $M=\gamma_{p}M_{H}$ and the approximation $g_{*}(T)\simeq g_{*,s}(T)$, we reproduce $\Omega_{\rm PBH}$ given in Ref. Carr:2009jm . We observe that the PBH energy fraction at the formation should be of order $\beta(M)\sim 10^{-9}$ so that the PBH relic abundance today does not exceed the required dark matter relic abundance. Thus the variance of the density contrast should be small and the probability density distribution should be concentrated so that the probability for the production of large density contrasts is exponentially suppressed in the tail of the probability density distribution.
5.2 Inflation-induced density perturbation
@@
Let’s first briefly review the PBH production in the inflation scenario. For the density fluctuations from the curvature perturbations during inflation, the variance can be estimated as
$\sigma^{2}=\int_{0}^{\infty}W^{2}(kR)\mathcal{P}_{\delta}(k)\frac{dk}{k},$ (41)
where $W(kR)$ is a volume-normalized Fourier transform of the window function smoothed over a comoving scale $R$ and can be assumed as Gaussian. On comoving hypersurfaces, the density perturbation can be related to the curvature perturbation by Green:2004wb
$\mathcal{P}_{\delta}(k)=\frac{4(1+\omega)^{2}}{(5+3\omega)^{2}}\left(\frac{k}{aH}\right)^{2}\mathcal{P}_{\mathcal{R}_{c}}(k).$ (42)
@@
$\mathcal{P}_{\mathcal{R}_{c}}(k)=A_{\mathcal{R}_{c}}\left(\frac{k}{k_{*}}\right)^{n_{s}-1},$ (43)
where the amplitude $A_{\mathcal{R}_{c}}$ is a constant that depends on the inflation physics, and $k_{*}$ is the reference scale. Current CMB and LSS observations on large scales ($k\lesssim{\rm Mpc^{-1}}$) have restricted the primordial curvature power spectrum at the order around $10^{-9}$. To have a sizable PBH formation for LIGO observation, the primordial curvature power spectrum should be of order $\mathcal{P}_{\mathcal{R}_{c}}(k)\sim 10^{-2}-10^{-1}$ (corresponding to a variance in the range $10^{-4}\lesssim\sigma^{2}\lesssim 10^{-2}$ which gives $\beta\sim 10^{-9}-10^{-6}$, see Fig. 10 in appendix D), which requires a blue-tilted spectral index $n_{s}\simeq 1.85$ at scale $k\sim 10^{6}\leavevmode\nobreak\ \rm Mpc^{-1}$ (corresponding to a PBH mass $M\sim 30M_{\odot}$asaki:2018dmp ; Byrnes:2018clq . On the other hand, the CMB observations have constrained the spectral index to be red-tilted, $n_{s}=0.968\pm 0.006$ at the scale $k_{*}=0.05\leavevmode\nobreak\ \rm Mpc^{-1}$. The relatively large blue-tilted spectral index at a small scale cannot be realized in the single-field slow-roll inflation but may be achieved in the running mass inflation or multi-scalar inflation models, in which the scale-dependent spectral index can be realized (see Refs. Stewart:1997wg ; Drees:2011hb ; Alabidi:2012ex ; Silk:1986vc ; Randall:1995dj ; GarciaBellido:1996mdl ; Clesse:2015wea for example).
5.3 Domain wall induced density perturbation
Compared with the inflation during which the density fluctuations are only induced at the surface of the last scattering, the metric perturbations are continuously generated by the domain wall network in the scaling regime. The domain wall tends to stretch in size and concatenate the nearby walls to form larger domains. This results in a domain wall network with large wall sizes and a small wall number at the late evolution stage. Therefore, one expects that the density perturbation and, thus, the power spectrum induced by the domain walls are most significant at the later stage. Such a picture is confirmed by recent simulations of the domain wall evolution to calculate the power spectrum Lazanu:2015fua . A simple estimate indicates that the power spectrum of the domain walls should have $\mathcal{P}_{w}\sim(G\sigma_{w}t)^{2}$, which is consistent with the variance given by Eq. (17). Thus, the density perturbations induced by the few-body system consist of domain walls that grow with time even in the radiation-dominated universe.
As has been shown in section 2.2 as well as the simulations, the energy density of the domain wall in the scaling regime is given by
$\rho_{w}=\frac{\mathcal{A}\sigma_{w}}{t}=2\mathcal{A}H\sigma_{w},$ (44)
where $\mathcal{A}$ depends on the number of walls. For $N\leq 6$ we take the values of $\mathcal{A}$ from the simulations Kawasaki:2014sqa , and for $N>6$ we take $\mathcal{A}=0.4N$. Therefore, the energy fraction of the wall is
$f_{w}=\frac{\rho_{w}}{\rho_{\rm rad}}=\frac{16\pi G\mathcal{A}\sigma_{w}}{3H}=\frac{32\pi G\mathcal{A}\sigma_{w}t}{3}.$ (45)

### v2

@@ 5 Critical collapse
Let us assume that the perturbed region is spherically symmetric and has an initial radius $R(t)$ larger than the Hubble scale. The locally overdense region can then constitute a spatially closed Universe with a metric given by [58, 59]
$ds^{2}=-dt^{\prime 2}+R^{2}\left(t^{\prime}\right)\left[\frac{dr^{2}}{1-\kappa r^{2}}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\varphi^{2}\right)\right],$ (5.1)
@@
$\rho=\frac{3}{8\pi G}\left(H^{2}+\frac{\kappa}{R^{2}}\right).$ (5.2)
The initial expansion rates of the background and the perturbed region can be set to be the same by choosing the coordinates [60, 59] so that the initial density contrast is given by
$\delta_{i}=\frac{\rho_{i}-\bar{\rho}_{i}}{\bar{\rho}_{i}}=\frac{3\kappa}{8\pi G\bar{\rho}_{i}R_{i}^{2}}=\frac{\kappa}{H_{i}^{2}R_{i}^{2}},$ (5.3)
@@
$\rho=\rho_{i}\left(\frac{R}{R_{i}}\right)^{-3(1+\omega)},$ (5.5)
@@
$\frac{\rho_{c}}{\rho_{i}}=\frac{\kappa}{R_{i}^{2}H_{i}^{2}+\kappa}\frac{R_{i}^{2}}{R_{c}^{2}}=\frac{\delta_{i}}{1+\delta_{i}}\frac{R_{i}^{2}}{R_{c}^{2}}.$ (5.6)
Together with Eqs. (5.5) and (5.6), we obtain the critical scale factor that the overdense region stops expanding
$\frac{R_{c}}{R_{i}}=\left(\frac{1+\delta_{i}}{\delta_{i}}\right)^{\frac{1}{1+3\omega}}.$ (5.7)
@@
$\frac{t_{c}}{t_{i}}=\left(\frac{1+\delta_{i}}{\delta_{i}}\right)^{\frac{3(1+\omega)}{2(1+3\omega)}}.$ (5.8)
The overdense region could contrast further against the pressure gradient if the self-gravitation $\Phi$ of the mass within the critical radius $R_{c}$ is larger than the internal energy $U$ of the overdense region, which gives [33]
$\Phi+U\sim-G\rho^{2}R_{c}^{5}+\omega\rho R_{c}^{3}\gtrsim 0.$ (5.9)
@@
$R_{c}\gtrsim R_{J}\simeq c_{s}t_{c},$ (5.10)
where $R_{J}$ is the Jeans scale and $c_{s}\simeq\sqrt{\omega}$. With Eqs. (5.7), (5.8), and (5.10), one finds the critical density criterion for PBH formation in the radiation-dominated Universe:
$\delta_{i}\gtrsim\delta_{c}\simeq c_{s}^{2}\simeq\frac{1}{3}.$ (5.11)
@@
$\frac{1}{3}\lesssim\delta_{i}\lesssim 1.$ (5.12)
The Hawking and Carr arguments for PBH production can be succinctly represented by the Newton potential criterion
$|\Phi|=\frac{3}{2}\left(\frac{k}{\mathcal{H}}\right)^{2}\delta\gtrsim 1,$ (5.13)
where we have omitted the subscript $N$ in Eq. (4.33). During the horizon crossing of mode $k$ with $k\simeq\mathcal{H}$, the Newton potential criterion equivalently requires $\delta\gtrsim\delta_{c}\simeq 0.5$. Using the Poisson distribution in Eq. (3.1) with the variance in Eq. (3.18), the ensemble-averaged density fluctuation from the isocurvature perturbation is given by
$\bar{\delta}(t_{\rm ann})=\int_{0}^{\infty}\delta P(\delta)d\delta=\frac{f_{w}(t_{\rm ann})}{\sqrt{2\pi N}},$ (5.14)
where $t_{e}=t_{f}\simeq t_{\rm ann}$ has been used. Comparing this with Eq. (4.32), we observe that the mean initial isocurvature perturbation of the superhorizon mode is
$\bar{S}(0)=\frac{3}{\sqrt{2\pi N}}.$ (5.15)
For a network containing $N\lesssim 10$ domain walls, the isocurvature perturbation from the domain wall is $S(0)\sim\mathcal{O}(1)$. From Eq. (5.14), we observe that the density perturbation could exceed the critical value and form a PBH if the energy density fraction of the domain wall $f_{w}(t_{e})\sim\mathcal{O}(0.1)$ during the horizon crossing of the domain wall.
@@ 6.1 PBH relic abundance
$\beta(t_{f})=\frac{\rho_{\mathrm{PBH}}(t_{f})}{\rho_{\rm tot}(t_{f})}=\Omega_{\mathrm{PBH}}\left(\frac{H_{0}}{H_{f}}\right)^{2}\left(\frac{a_{0}}{a_{f}}\right)^{3},$ (6.1)
where $\Omega_{\mathrm{PBH}}\equiv\rho_{\mathrm{PBH}}\left(t_{0}\right)/\rho_{\text{crit}}$ (with $\rho_{\rm crit}=3H_{0}^{2}/(8\pi G)$) is the density of the PBH relic abundance at present and 0 denotes the present time. In the second equality of Eq. (6.1) we have used $\rho_{\rm PBH}(t_{f})=\rho_{\rm PBH}(t_{0})(a_{0}/a_{f})^{3}$. It is commonly related to the dark matter relic abundance by $\Omega_{\mathrm{PBH}}=f_{\mathrm{PBH}}\Omega_{\mathrm{CDM}}$, with the fraction of dark matter $f_{\rm PBH}$ in the form of PBHs.
It is expected that the PBHs formed from the Hubble horizon (re-)entry of the large density perturbations could have masses at the same order as the associated horizon mass $M_{H}\simeq 2\times 10^{15}(t/1s)M_{\odot}$. To further estimate the PBH production, let’s take the naive relation between the PBH mass $M$ and the horizon mass from critical phenomena [37, 38]
$M=\gamma_{p}M_{H},$ (6.2)
@@
$M_{H}=\frac{4\pi}{3}\rho_{\rm rad}r_{H}^{3}=\frac{1}{2GH}=\frac{3\sqrt{5}M_{\rm pl}^{3}}{4\pi^{3/2}g_{*}^{1/2}(T)T^{2}}.$ (6.4)
Taking advantage of the Press-Schechter formalism, the PBH energy fraction can be expressed as [61]
$\beta(M)=2\int_{\delta_{c}}^{\infty}\frac{M}{M_{H}}P(\delta)d\delta,$ (6.5)
where we consider a Gaussian probability distribution (3.19) for the density contrast. Note that we use an upper limit of infinity for the integration and confirm that it is not different from the integration result with an upper limit of $\delta=1$ from the requirement of avoiding separate Universes [33, 59]. A recent work [62] argued that this upper limit was a result of the geometry of the selected slicing rather than as a consequence of avoiding separate Universes, which instead led to a divergence of density fluctuations on the superhorizon scale. With the PBH mass given by Eq. (6.2), the fraction of PBHs at the formation time is
$\beta(M)=\gamma_{p}{\rm erfc}\left(\frac{\delta_{c}}{\sqrt{2}\sigma}\right),$ (6.6)
where ${\rm erfc}(x)$ is the complementary error function (see appendix D for more details). The critical density contrast $\delta_{c}$ varies with the equation of state $\omega$. For radiation phase $\delta_{c}=0.453$. We determine $\delta_{c}$ with the $\delta_{c}-\omega$ relation revealed by the simulations in Ref. [63] (see section 7).
A more precise PBH mass formation was obtained from the numerical computations on the collapse of a relativistic fluid, which indicated a simple scaling law for the PBH mass distribution function near the critical threshold [64, 65, 66]
$M=CM_{H}(\delta-\delta_{c})^{b},$ (6.7)
where the constant parameter $C=3.3$ and $b=0.36$ for the formation of PBH during radiation domination [67]. Note that these parameters may be affected by the phase of cosmic matter. Then the PBH energy fraction is given by
$\beta(M)=2\int_{\delta_{c}}^{\infty}C(\delta-\delta_{c})^{b}P(\delta)d\delta.$ (6.8)
In this work, we use the PBH mass (6.2) by default, unless otherwise stated.
@@
$s_{f}a_{f}^{3}=s_{0}a_{0}^{3},\leavevmode\nobreak\ \leavevmode\nobreak\ {\rm with}\leavevmode\nobreak\ \leavevmode\nobreak\ s(T)=\frac{2\pi^{2}}{45}g_{*,s}T^{3},$ (6.9)
where the current Universe entropy density $s_{0}=2891.2\leavevmode\nobreak\ \rm cm^{-3}$. Together with Eqs. (6.1)-(6.3) we obtain the the current PBH relic density abundance
$\Omega_{\mathrm{PBH}}$ $=\frac{2\pi G\beta(M)s_{0}g_{*}(T)T}{H_{0}^{2}g_{*,s}(T)}=(3\sqrt{5\pi})^{1/2}\frac{\beta(M)s_{0}}{H_{0}^{2}(M_{\rm pl}M_{H})^{1/2}}\frac{g_{*}^{3/4}(T)}{g_{*,s}(T)}$ (6.10)
@@
where the Hubble expansion rate is $H_{0}=100h\leavevmode\nobreak\ \mathrm{km\leavevmode\nobreak\ s}^{-1}\leavevmode\nobreak\ \mathrm{Mpc}^{-1}$ with $h=0.68$ from the Planck observation [68]. With the relation $M=\gamma_{p}M_{H}$ and the approximation $g_{*}(T)\simeq g_{*,s}(T)$, we reproduce $\Omega_{\rm PBH}$ given in Ref. [69]. We observe that the PBH energy fraction at the formation should be of order $\beta(M)\sim 10^{-9}$ so that the PBH relic abundance today does not exceed the required dark matter relic abundance. Thus the variance of the density contrast should be small and the probability density distribution should be concentrated so that the probability for the production of large density contrasts is exponentially suppressed in the tail of the probability density distribution.
6.2 Density fluctuation from inflation-induced curvature perturbation
@@
Let’s first briefly review the PBH production in the inflation scenario. For the density fluctuations from the curvature perturbations during inflation, the variance is in general estimated by Eq. (3.17). On comoving hypersurfaces, the density perturbation can be related to the curvature perturbation by [70]
$\mathcal{P}_{\delta}(k)=\frac{4(1+\omega)^{2}}{(5+3\omega)^{2}}\left(\frac{k}{aH}\right)^{2}\mathcal{P}_{\mathcal{R}_{c}}(k).$ (6.11)
@@
$\mathcal{P}_{\mathcal{R}_{c}}(k)=A_{\mathcal{R}_{c}}\left(\frac{k}{k_{*}}\right)^{n_{s}-1},$ (6.12)
where the amplitude $A_{\mathcal{R}_{c}}$ is a constant that depends on the inflation physics, and $k_{*}$ is the reference scale. Current CMB and LSS observations on large scales ($k\lesssim{\rm Mpc^{-1}}$) have restricted the primordial curvature power spectrum at the order around $10^{-9}$. To have a sizable PBH formation for LIGO observation, the primordial curvature power spectrum should be of order $\mathcal{P}_{\mathcal{R}_{c}}(k)\sim 10^{-2}-10^{-1}$ (corresponding to a variance in the range $10^{-4}\lesssim\sigma^{2}\lesssim 10^{-2}$ which gives $\beta\sim 10^{-9}-10^{-6}$, see Fig. 11 in appendix D), which requires a blue-tilted spectral index $n_{s}\simeq 1.85$ at scale $k\sim 10^{6}\leavevmode\nobreak\ \rm Mpc^{-1}$ (corresponding to a PBH mass $M\sim 30M_{\odot}$[37, 67]. On the other hand, the CMB observations have constrained the spectral index to be red-tilted, $n_{s}=0.968\pm 0.006$ at the scale $k_{*}=0.05\leavevmode\nobreak\ \rm Mpc^{-1}$. The relatively large blue-tilted spectral index at a small scale cannot be realized in the single-field slow-roll inflation but may be achieved in the running mass inflation or multi-scalar inflation models, in which the scale-dependent spectral index can be realized (see Refs. [71, 72, 73, 74, 35, 75, 76] for example).
6.3 Density fluctuation from isocurvature perturbation induced by domain wall network
As demonstrated in section 2.2, the isocurvature perturbations from the domain wall network are produced during the horizon crossing of the domain wall. We consider the PBHs due to the density fluctuation arising from the isocurvature perturbation at the horizon crossing time of the domain wall. At this stage, the bias potential becomes significant, which could lead to the annihilation of the domain wall network. Therefore, we have $t_{e}=t_{f}\simeq t_{\rm ann}$. The energy density of the domain wall is related to the wall’s scale by $\rho_{w}=\sigma_{w}/\bar{L}$. At the horizon crossing time $t_{e}$, $\bar{L}\simeq t/\mathcal{A}$, and the domain wall energy density is then given by
$\rho_{w}=\frac{\mathcal{A}\sigma_{w}}{t}=2\mathcal{A}H\sigma_{w},$ (6.13)
where $\mathcal{A}$ depends on the number of walls. For $N\leq 6$ we take the values of $\mathcal{A}$ from the simulations [77], and for $N>6$ we take $\mathcal{A}=0.4N$. Therefore, the energy fraction of the wall is
$f_{w}=\frac{\rho_{w}}{\rho_{\rm rad}}=\frac{16\pi G\mathcal{A}\sigma_{w}}{3H}=\frac{32\pi G\mathcal{A}\sigma_{w}t}{3}.$ (6.14)